OpenStudy (anonymous):

Explain if :

OpenStudy (anonymous):

\[x = 9 - 4 \sqrt{5}, find \ the \ value \ of \ \ x^2 + \frac { 1 }{ { x }^{ 2 } } \]

OpenStudy (anonymous):

@Callisto

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

help

OpenStudy (anonymous):

@\\@Hally canu

OpenStudy (anonymous):

@vbfn

OpenStudy (anonymous):

@vbfn help

OpenStudy (anonymous):

plzz help

OpenStudy (anonymous):

\[x=9-4\sqrt{5}\] then what is x^2?

OpenStudy (anonymous):

x^2=9*9-4*4*5 since sqrt5*squrt5=5

OpenStudy (anonymous):

x^2=81-90=9

OpenStudy (anonymous):

x^2+(1/x^2)= 9+(1/9) and I'm sure you can do that.

OpenStudy (anonymous):

thanks fr u r help

OpenStudy (anonymous):

Is the answer 82/9

OpenStudy (anonymous):

@vbfn

OpenStudy (anonymous):

yes!